OpenStudy (anonymous):

HELP IT'S REALLY EASY The profit depends on the amount x that you charge for a box. The equation p= -0.5x^2 + 25x - 150 models the profit. What's the smallest amount that can be charged to make a profit of at least $125?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@AGummiBear55

OpenStudy (anonymous):

@Joshyface

OpenStudy (anonymous):

CMON @Joshyface YOU WERE TYPING I HELPED YOU LOL please

OpenStudy (anonymous):

@AravindG @abb0t @Aylin @ashlynn143 @annahampton @brinethery @breja @binarymimic @BradNasty @CalebBeavers @Carter15 @christinewilliams @dydlf @dirtrider @dumbcow

OpenStudy (anonymous):

@Joshyface

OpenStudy (anonymous):

@BradNasty come on please

OpenStudy (anonymous):

THANK YOU. cmon someone help

OpenStudy (anonymous):

@jim_thompson5910 I KNOW UR ON

OpenStudy (anonymous):

@dydlf @etaylor13579 @Firejay5 @ganeshie8 @girl05 @GodBlessYou @henryrodriguez713 @Helicopter @inkyvoyd @idksucks @iJustAteCheese @Itel @JuLyBaBe97 @JetWave600 @jennaaanicoleee @karatechopper @k.rajabhishek

OpenStudy (inkyvoyd):

p= -0.5x^2 + 25x - 150 seek smallest value of x such that p>=125 does that make sense?

OpenStudy (anonymous):

noo :(

OpenStudy (inkyvoyd):

You want -0.5x^2 + 25x - 150 >=125. Solve the quadratic inequality and eliminate extraneous solutions.

OpenStudy (anonymous):

lemme try and do it WAIT HERE lol

OpenStudy (firejay5):

16 = $122

OpenStudy (firejay5):

I am not for sure if it's right though

OpenStudy (anonymous):

okay so did you factor it?

OpenStudy (firejay5):

me, no I type it into my graphing calculator and that's what I got

OpenStudy (inkyvoyd):

0.5x^2 + 25x - 150 >=125 x^2+50x-300>=250 (x^2+50x+625)>=550+625 (x+25)^2>=1175

OpenStudy (anonymous):

and you did it like this is it correct?

OpenStudy (inkyvoyd):

it might be it might not be. That's what you have to judge.

OpenStudy (anonymous):

so What's the smallest amount that can be charged to make a profit of at least $125

OpenStudy (inkyvoyd):

You asked me the answer? Don't do that. It's 235.2

OpenStudy (anonymous):

so its wrong?

OpenStudy (inkyvoyd):

Figure out the answer for yourself.

OpenStudy (anonymous):

I don't know what to do though.

OpenStudy (inkyvoyd):

p= -0.5x^2 + 25x - 150 seek smallest value of x such that p>=125 does that make sense? you said that doesn't make sense - what part of it doesn't make sense?

OpenStudy (anonymous):

nevermind i got it